differential equations

Use the Laplace transform to solve the given initial-value problem. y - y = e cost, y(0) = 0, y(O) = 0 y(t) =
Answer #1

ه - (o) : (مك - - y = et cost Laplace transforms s? Y(S)- SYCS) = S- ) | + را- د) ۸( ۲) ۔ رو 1 +(1-ES (سری )s S - ۸ (۲) ۔2 - کبط ] : (ر) د 2. s 2 2 + 2 -او اس S سے لا = (ر) 7 (د S ا أ - ي) 2. 12-اراک sint tet تا دم اے - | cos 21 = y(t) = 2.